OpenStudy (anonymous):

If r = -6, s = -2, and t = -3 then -t(r - s) = A) -24 B) -12 C) 12 D) 24

OpenStudy (anonymous):

-12

OpenStudy (rane):

subsitude them in the equation

OpenStudy (anonymous):

kk thanks and i did

OpenStudy (rane):

-t(r-s) = = -3(-6 -2) = -3 * -8 = 24

OpenStudy (anonymous):

Uh, no, it equals -12 @RANE

OpenStudy (rane):

its not

OpenStudy (rane):

look at the working out, then tell me where am i wrong

OpenStudy (rane):

if u dont belive me, use calculator to find out

OpenStudy (anonymous):

http://www.wolframalpha.com/ put the equation in there LOL.

OpenStudy (anonymous):

it also becomes a POSITIVE 3, not a negative. two negatives make a positive

OpenStudy (rane):

ummm. i think u haven't checked it properly....this web also gives 24 as a ans

OpenStudy (anonymous):

so which shud i pick..etha 24 or -12

OpenStudy (rane):

use calculator to make sure your answer @Pookiieluvzyhu

OpenStudy (rane):

or also the web recommended by @whalexnuker

OpenStudy (anonymous):

-12

OpenStudy (rane):

wht the.... did u subsitude the numbers in the equations properly

OpenStudy (anonymous):

@RANE put "if r = -6, s = -2, and t = -3 then -t(r - s) =" into that site... it gives you -12. Stop trying to substitute it on your own.

OpenStudy (rane):

can i see how did you subsitude in the formula

OpenStudy (rane):

your new equation

OpenStudy (rane):

@Pookiieluvzyhu did u check the ans by calc

OpenStudy (anonymous):

yes its -12...thank 4 all your help though

OpenStudy (rane):

wht!!

OpenStudy (anonymous):

@RANE :]

OpenStudy (anonymous):

sorry rane..but thatz what i got..

OpenStudy (rane):

as u wish bt i'm still saying that -12 is wrong

OpenStudy (ivettef365):

it is -12

OpenStudy (ivettef365):

What you did wrong is that the equation should be 3(-6+2)

OpenStudy (rane):

s = -2, not +2

OpenStudy (rane):

read the q'

OpenStudy (ivettef365):

yes but you have the following: -t(r-s) -(-3) (-6-(-2))

OpenStudy (ivettef365):

Therefore: 3(-6+2)

OpenStudy (rane):

oooooo....my bad i didn't see the -ve sign with the t

OpenStudy (rane):

sorrryyyy guys

OpenStudy (anonymous):

lol its cool..y r stilll smart lol :]

OpenStudy (rane):

my bad

OpenStudy (rane):

@whalexnuker yeah u r right ..i just missed the -ve sign with the t